OpenStudy (anonymous):

What is the phase shift for the equation: y=-3sec(π-2x)+5

OpenStudy (anonymous):

@pooja195 @Hero @e.mccormick @SolomonZelman

OpenStudy (lynfran):

this equation can be written as y=5-3sec(-2x+pi) where the phase shift would be -2x+pi=0

OpenStudy (lynfran):

solve for x ...and thats ur phase shift

OpenStudy (lynfran):

what do u get?

OpenStudy (anonymous):

2x/pi

OpenStudy (lynfran):

no -2x+pi=0 -2x=-pi divide by -2 x=pi/2

OpenStudy (anonymous):

so the phase shift is pi/2

OpenStudy (lynfran):

yep

OpenStudy (anonymous):

wait I have one more question, is there a such thing of having a phase shift of 0 or is that considered not a phase shift @LynFran

OpenStudy (lynfran):

not a phase shift .....a phase shift mean that there a vertical shift ...either to the right or left of the graph

OpenStudy (anonymous):

\[y=-2\sin(5/4x) \]

OpenStudy (lynfran):

and that new phase shift determines where the graph starts

OpenStudy (anonymous):

^Does that equation have no phase shift then^

OpenStudy (lynfran):

yes that equation has no phase shift

OpenStudy (lynfran):

it just have the amplitude which is |-2|=2 and its period which is 2pi divided by 5/4x

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

wait I thought the amplitude would just be -2

OpenStudy (anonymous):

the amplitude does not have absolute value

OpenStudy (anonymous):

F(t) = Af(Bt – C) + D, A is the ampitude

OpenStudy (lynfran):

no amplitudes are always the absolute value of the # infront the function /curve .... which means its always positive...

OpenStudy (anonymous):

i think that you are wrong

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

OH WAIT YOU ARE RIGHT, I WAS THINKING ABOUT PERIODICITY

OpenStudy (anonymous):

IM SO SORRY

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

its ok